# Self-Controlled Diffusion for Denoising in Scientific Imaging

### v1

@@ 4.3 Prediction
Conceptually, the prediction Algorithm 3, depicted in Figure 4, is similar to the improved sampling algorithm from [19], with two key novel features arising from the use of the two regression values. First, we estimate the initial diffusion step based on a dry-run of the neural network. In this step, the input timestep is set to $T$ the total number of timesteps used during training.
@@
Moreover, given the definition of $\hat{r}_{t_{x}}$, its value approaching zero can serve as a useful indicator to flag potential hallucinations or unreliable predictions. This is particularly valuable in ensuring the robustness of the denoising process and detecting situations where the model might be overfitting or making implausible predictions. Usage of $\hat{r}_{t_{x}}$ allows getting the best denoised image, based on the evaluation, rather then conventionally selected image from the last denoising step.
@@
Note that the proposed prediction schedule has no reliance on clean images (Master Patterns), which is crucial for the application.
@@ 4.4 Architecture and Implementation details
All experiments used $128\times 128$ images and a standard UNet architecture with two residual layers per block, following [5, 19]. A key deviation was downscaling only in the first encoder block. The auxiliary head had two DownBlocks (128 and 64 filters), followed by GlobalAveragePooling, batch normalization, a fully connected layer with 64 neurons, and an output layer with two regression outputs. Diffusion step values were encoded with a sine-cosine positional embedding and processed through two fully connected layers with SiLU activation.
We used the Huber loss function (Eq. 4) with $\delta=0.05$ for both tasks, trained with the AdamW optimizer [13], starting with a learning rate of $10^{-5}$ for 256 epochs. The learning rate was halved after 18 epochs without improvement in SSIM (denoising) or $L_{2}$ (regression) on validation.
Two augmentations were applied: MixUp [6] with a probability of 0.25 and a maximum added pattern weight of 0.25, and 10% rate of Gaussian noise-only images paired with zero targets during the denoising phase. $s_{\text{noise}}$ and $p_{\text{noise}}$ were set to 0.5 during the experiments.
Normalization steps were omitted for brevity in Algorithms 2 and 3, but $x_{t}$ values were normalized to zero mean and unit variance. The decoder’s final activation was tanh, scaled by 3.5, while the quality prediction head used a sigmoid activation. $T=384$ in all experiments.
@@
$\mathcal{L}(\hat{x},x)=\begin{cases}0.5(\hat{x}-x)^{2},&\text{if }|\hat{x}-x|\leq\delta,\\ \delta\left(|\hat{x}-x|-0.5\delta\right),&\text{otherwise.}\end{cases}$ (4)
Figure 5: The results of image restoration. N - the number of restoration cycles (the maximum number is $T=384$). Step is the model’s diffusion step estimation (the lower the more restored the image is according to model estimation). Initial quality (estimated quality) is the predicted quality of the original pattern $\hat{q}$; denoising quality - the model estimation of cross-correlation result between restored image and the ground truth $\hat{r}$. The a) and b) - EBSD patterns for Nickel sample of different quality; c) - instrumental noise sample.
@@ 5 Results and discussion
In this section, we evaluate the impact of different components of our approach on the denoising performance. Specifically, we assess the contributions of various training and inference strategies used in our diffusion-based denoising framework.
@@
To ensure a fair comparison, all experiments are conducted using the same fixed subset of the dataset, with identical hyperparameters, augmentation strategies, and learning rate schedules. The evaluation is performed on a hold-out test set, where we measure image quality using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM).
@@
Since our model is trained to predict a denoised image directly from any diffusion step, we first compare different training strategies for the denoising component. This evaluation is based on the model’s ability to directly reconstruct the clean image from a noisy input. Table 1 summarizes the results obtained using different methods for generating noisy training samples, $x_{t}$.
Process PSNR ($\uparrow$) SSIM ($\uparrow$)
Raw data 14.1317 0.5235
Baseline [19] 15.7447 0.6632
Our method (Algorithm 1) 17.6460 0.7728
Table 1: Ablation study of different training procedures for diffusion-based denoising. Raw data represents the test dataset without denoising. Baseline corresponds to the conventional linear algorithm from [19]. Our method follows Algorithm 1, demonstrating significant improvements in both PSNR and SSIM.
The results show that our proposed training strategy significantly outperforms the baseline approach, achieving a noticeable increase in both PSNR and SSIM. This confirms the effectiveness of our training novelties: quality-defined control of training input image generation and linear extrapolation of experimental images towards more degraded state then the experimentally observed image.
Next, we investigate different inference-time denoising strategies, analyzing how modifications to the step selection process impact image quality. Table 2 presents the results of various step selection algorithms.
Process PSNR ($\uparrow$) SSIM ($\uparrow$)
Baseline [19] 17.6737 0.7742
Adjusted ($t_{x}=T\cdot(1-q)$) 17.6771 0.7760
Our method 17.7964 0.7791
Our method (best-step selected) 18.0589 0.7890
Table 2: Ablation study of different denoising procedures. Baseline represents the standard linear denoising approach from [19]. Adjusted modifies the baseline by determining the starting diffusion step based on predicted pattern quality, as $t_{x}=T\cdot(1-\hat{q})$. Our method further improves on this by incorporating dynamic selection of the next diffusion step based on predicted $\hat{r}$, where $t_{\text{next}}=T\cdot(1-\hat{r})$. This method follows the entire Algorithm 3 except for the best-step selection (last line of the Algorithm 3. Our method (best-step selected) refines the previous approach by selecting the final pattern as the one corresponding to the highest predicted $\hat{r}$.
The results indicate that dynamically adjusting the diffusion starting step based on predicted quality leads to consistent performance improvements. The best performance is achieved when the final denoised pattern is selected based on the highest predicted $\hat{r}$, demonstrating the utility of assessing denoising progress in an adaptive manner and is an evidence of the effectiveness of our denoising progress assessment. This result highlights that for some patterns, optimal denoising occurs at an intermediate step, with further iterations potentially degrading quality. The ability to adaptively stop at the best step offers a significant advantage over fixed-step approaches, highlighting the importance of an adaptive stopping criterion for achieving optimal reconstruction while avoiding unnecessary processing.
Figure 6: Predicted values of $\hat{q}$ (pattern quality) and $\hat{r}_{denoise}$ (denoised quality) for a proper dataset, including actual EBSD patterns and images with instrumental noise. The results demonstrate that for real EBSD patterns, the predicted denoising progress values are significantly higher, whereas for noise data, the predicted values remain low for the denoising process.
@@
The observed results in Figure 6 reveal a clear distinction between actual images and instrumental noise samples, based on the predicted values of $\hat{q}$ and $\hat{r}_{denoise}$. For real EBSD patterns, the final predicted denoising progress values are significantly higher, indicating the model’s ability to recognize meaningful patterns. In contrast, for noise data, the predicted values remain consistently low, reflecting the absence of structure in the input data. These findings suggest that the predicted values can serve as a criterion for differentiating between actual images and noise, offering a reliable approach for noise detection and filtering in practical applications.
@@ 6 Conclusion
We propose a dynamic method for selecting diffusion steps based on predicted pattern quality, $\hat{q}$, and denoised quality, $\hat{r}_{denoise}$. This adaptive strategy ensures optimal termination of the denoising process, integrates quality-defined control for generating noisy training samples, and extrapolating experimental images to more degraded states, our approach significantly improves denoising performance compared to baseline methods. The predicted values of $\hat{q}$ and $\hat{r}_{denoise}$ also enable noise-only data detection and filtering. These contributions enhance the denoising process, improving reconstruction quality and robustness.
This method could prove useful in future EBSD measurements, allowing for reduced exposure times while sacrificing some pattern quality, and subsequently restoring the information during post-processing. Furthermore, the suggested approach lessens the reliance on the experimentalist’s selection of the image processing sequence and could be useful beyond EBSD context.

### v2

@@ 4.3 Prediction
Conceptually, the prediction Algorithm 3, depicted in Figure 3, is similar to the improved sampling algorithm from [27], with two key novel features arising from the use of the two regression values. First, we estimate the initial diffusion step based on a dry-run of the neural network (the input timestep is set to $T$ during the run).
@@
Moreover, given the definition of $\hat{r}_{t_{x}}$, its value approaching zero can serve as a useful indicator to flag potential hallucinations or unreliable predictions. This is particularly valuable in ensuring the robustness of the denoising process and detecting situations where the model might be making implausible predictions. Usage of $\hat{r}_{t_{x}}$ allows getting the best denoised image, based on the evaluation, rather then conventionally selected image from the last denoising step.
@@
Note that the proposed prediction schedule has no reliance on clean images (Master Patterns).
@@ 4.4 Architecture and Implementation details
All experiments used $128\times 128$ images and a UNet style CascadedGaze architecture [10], following the Real Image Denoising protocol from [10] with a network width of 60. The encoder contains [2, 2, 4, 6] CascadedGaze blocks, the middle layer has 10 NAF blocks [5], and the decoder comprises [2, 2, 2, 2] blocks. The auxiliary head is composed of an InstanceNorm layer, followed by two convolutional blocks that each downsample the input by a factor of 2 and incorporate a skip connection. After global‑average pooling, the resulting features are fed into a 3‑layer MLP that employs batch normalization and a dropout with rate of 0.1. The hidden layer sizes are 256 and 128, and the final layer outputs two regression values, which are passed through a sigmoid activation. Diffusion‑step values are encoded with 128‑dimensional sine–cosine positional embeddings and then projected through a fully connected layer. The SiLU function is used as the non‑linearity in this network.
We used the L1 loss function for both tasks, trained with the AdamW optimizer [20] with $\beta_{1}=0.9,\beta_{2}=0.9$, starting with a learning rate of $10^{-3}$ and decays to $10^{-7}$ using a cosine annealing scheduler over 96 epochs.
Two augmentations were applied during the first stage training: MixUp [28] with a probability of 0.25 and a maximum added pattern weight of 0.25, and 10% rate of Gaussian noise-only images paired with zero targets. $s_{\text{noise}}$ and $p_{\text{noise}}$ were set to 0.5. In Algorithms 2 and 3 $x_{t}$ values were normalized to range [0, 1]. $T=384$ in all experiments.
@@ 5 Results and discussion
In this section, we are comparing different training and inference strategies and evaluating the impact of our framework’s components on denoising performance.
Figure 4: Visual qualitative comparison of different EBSD patterns restoration methods.
@@
All experiments use the same fixed data subsets, hyperparameters, and augmentations, with evaluation on a hold-out test set using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM).
@@
Since our model can predict a denoised image from any diffusion step, we first compare training strategies for the denoising component. As EBSD denoising has received little attention, we benchmark against representative scientific‑image denoising methods (Table 1), measuring how well a clean image can be reconstructed directly from noisy input.
One step process PSNR ($\uparrow$) SSIM ($\uparrow$)
Raw data 17.2198 0.5160
BM3D [9, 21] 18.5254 0.6310
NAFNet [5, 26] 24.9287 0.7991
CascadedGaze [10] 24.9309 0.7931
DIFFRACT 25.0333 0.8023
DIFFRACT with $t_{x}$ assessment 25.0337 0.8067
Table 1: Comparison of different procedures for direct patterns restoration. Raw data represents the test dataset without denoising. BM3D, NAFNet and CascadedGaze are reference training-free and DL methods. Our method is similar to the conventional linear algorithm from [27], expanded by incorporation of input data quality and linear extrapolation to more degraded images. Our method with $t_{x}$ assessment follows one step prediction of $\hat{x}_{0|t}$ by Algorithm 3 (lines 1-5).
Diffusion process PSNR ($\uparrow$) SSIM ($\uparrow$)
DDPM [11] 25.1763 0.8043
DDPM-NLC [1] 25.2081 0.8132
Baseline 25.0620 0.8001
[…]
+ best step selection, based on $\hat{r}_{t_{x}}$ 25.2825 0.8093
Table 2: The first two rows report the performance of the reference DDPM and DDPM‑NLC models, both applied with deterministic sampling ($\eta=0$). Baseline represents the standard linear denoising approach similar to [27]. The remaining rows aprogressively add the components of our proposed denoising scheme with the final values demonstrating results of the complete Algorithm 3.
BM3D [21, 9] which is a widely used collaborative filtering non-local image denoising algorithm, performs far below learning-based methods, indicating that our dataset contains noise and structures beyond classical capabilities, despite BM3D’s success on synthetic noise and other imaging tasks [22, 25]. Since our method utilizes the same base architecture as CascadedGaze [10], which represents a state-of-the-art evolution of the widely used NAFNet convolutional architecture employed in scientific image denoising [26, 5], we trained the models with the original training protocol as a reference baseline. The results show that our training strategy outperforms the baseline approach, achieving an increase in both PSNR and SSIM metrics, confirming the value of quality-defined input generation and linear extrapolation to more degraded states. Predicted image quality for $t_{x}$ selection benefits even direct denoising without diffusion.
Figure 5: a) IPF maps of the original and restored datasets using DIFFRACT, with patterns shown before any image processing or neighboring-pattern averaging. b) Confidence Index (CI) values distribution for IPF map of the original (blue) and restored (orange) patterns.
Next, we investigate different inference-time diffusion-based denoising strategies, analyzing how modifications to the step selection process impact image quality (Table 2). The results demonstrate that dynamically adjusting the diffusion starting step based on predicted quality improves the performance of iterative denoising as well. The approach of dynamically selecting the next diffusion step based on the assessment of the denoising progress $\hat{r}$ proves beneficial and adds to the metric values. Selecting the final denoised pattern using the highest predicted $\hat{r}$ demonstrates the effectiveness of adaptive denoising progress assessment. This result shows that optimal denoising can occur at intermediate steps, where further iterations may actually degrade quality. Adaptive stopping at the optimal step provides a clear advantage over fixed-step methods, emphasizing the importance of quality-based adaptive criteria for achieving optimal reconstruction. Examples of iterative denoising results are presented in Fig. 4 and visual comparison of different methods in Fig. S7.
We also compare our method with other diffusion‑scheduling approaches (Table 2). For these experiments, we used the same architecture as our model and the same data preparation, including artificial image degradation and augmentations, with the only difference being the scheduling. In particular, we benchmark against restoration with DDPM and DDPM‑NLC (Algorithm 1 from [1] with $\eta=0$). The former is a base diffusion approach, whereas the latter is the closest to ours: its adaptive schedule varies the amount of noise at each step, whereas our approach selects the first and the next step number dynamically. Although the DDPM‑NLC approach achieves slightly better SSIM value, it does not provide a straightforward way to stop denoising at a preferred image‑quality level (the relative pattern quality should be preserved to maintain surface contrast differences like grain boundaries in the results after indexing) or to assess raw‑image quality to reduce the risk of hallucination.
@@
The observed results in Fig. S8 reveal a clear distinction between actual images and instrumental noise samples, based on the predicted values of $\hat{q}$ and $\hat{r}_{denoise}$. For real EBSD patterns, the final predicted denoising progress values are significantly higher, indicating the model’s ability to recognize meaningful patterns. In contrast, for noise data, the predicted values remain consistently low, reflecting the absence of structure in the input data. These findings suggest that the predicted values can serve as a criterion for differentiating between actual images and noise, offering a reliable approach for noise detection and filtering in practical applications.
Fig. 5 shows IPF maps of Nickel sample measured with extremely low exposure time - 1 ms. Each point’s color and brightness result from EBSD pattern indexing with SI. In the “original” dataset, point 2 is brighter than point 1 due to a weak Kikuchi line in the EBSD pattern (indicated by white arrows) that increases the OIM9 software’s quality parameter Confidence Index (CI) - confidence in crystal orientation assignment (point 1 does not relieve any visible lines). The same line, restored, is visible in the processed pattern. Original dataset patterns were enhanced using classical image processing (Adaptive Histogram Equalization, background subtraction, and neighboring-pattern averaging), whereas restored patterns were processed only by averaging neighboring patterns. Group of pixels with the same hue represent material grains - their size distribution and crystal orientation are the main information extracted from the EBSD analysis. Notably, the restored dataset preserved the surface-relief texture, reflecting the quality-assessment feedback mechanism that maintains differences between initially higher- and lower-quality patterns. This is consistent with the CI histogram (Fig. 5-b), which shows a wider CI distribution for the restored patterns. Thinner dark lines between grains in the “restored” map indicate better indexing quality and more precise grain-size estimation. Moreover, ”restored” map reveal more smaller grains through improved indexing of restored patterns. At the end, our method preserves critical surface features and enhances grain boundary clarity, enabling more accurate grain-size estimation and revealing smaller grains, while requiring less preprocessing than classical approaches.
Our ablation study demonstrates that incorporating image quality assessment into the diffusion denoising process yields consistent improvements over baseline methods. The proposed quality-defined control and noisy pattern extrapolation strategies, combined with adaptive timestep selection, collectively improve denoising metrics. The limited performance of traditional methods like BM3D confirms that EBSD pattern complexity necessitates learning-based approaches, justifying the effectiveness of quality-aware training and adaptive denoising strategies for challenging real-world scenarios in scientific imaging.
@@ 6 Conclusion
We propose the DIFFRACT - a dynamic method for selecting diffusion steps based on predicted pattern quality, $\hat{q}$, and denoised quality, $\hat{r}_{denoise}$. This adaptive strategy: 1) ensures optimal termination of the denoising process; 2) integrates quality-defined control for generating noisy training samples; 3) extrapolating experimental images to more degraded states. To the best of our knowledge, this is the first time such an approach has been applied to EBSD pattern restoration. Our approach significantly improves denoising performance compared to baseline methods. The predicted values of $\hat{q}$ and $\hat{r}_{denoise}$ also enable noise-only data detection and filtering. These contributions enhance the denoising process, improving reconstruction quality and robustness.
This method is proved useful in EBSD measurements, allowing for reduced exposure times while sacrificing some pattern quality, and subsequently restoring the information during post-processing. Furthermore, the suggested approach lessens the reliance on the experimentalist’s selection of the image processing sequence and could be useful beyond EBSD context.
@@ 7 Acknowledgments
Supplementary Material
Figure S6: An example of a Master Pattern (MP) and its components identified during Spherical Indexing. The same components of the MP served as ground truth in model training.
Figure S7: The results of image restoration. NN — the number of restoration cycles (maximum $T=384$). Step is the model’s diffusion step estimation (lower values indicate a more restored image, according to the model). Initial quality (estimated quality) is the predicted quality of the original pattern, $\hat{q}$; denoising quality is the model’s estimate of the cross-correlation between the restored image and the ground truth, $\hat{r}$. a) and b) - EBSD patterns for Nickel samples of different quality; c) - instrumental noise sample.
Figure S8: Predicted values of $\hat{q}$ (pattern quality) and $\hat{r}_{\text{denoise}}$ (denoised quality) for a dataset including actual EBSD patterns and images with instrumental noise.